# I Predict Therefore I Am: Is Next Token Prediction Enough to Learn Human-Interpretable Concepts from Data?

### v2

@@ G.1 LLMs Mimic the Human World Model, Not the World Itself
Interestingly, our texts reflect our mental activities and emotions, providing a window into this compressed world model. A recent study reveals a striking disparity between the limited information throughput of human behavior (approximately $10$ bits/s) and the vast sensory input available (around $10^{9}$ bits/s) (Zheng & Meister, 2024). This suggests that the human world model is an efficient, compressed abstraction, enabling us to reason, predict, and make decisions effectively. Despite this compression, humans have flourished as the dominant species on Earth, demonstrating the power of such a streamlined model.
@@
G.2 LLMs Versus Pure Vision Models: A Fundamental Difference
@@
Appendix H More Results on Llama-3 and DeepSeek-R1
We conduct additional experiments on recent LLMs, including Llama-3 and DeepSeek-R1, to further evaluate our findings. The experimental setup strictly adheres to the settings described in Section F.2. This enables a comprehensive investigation of our findings, ensuring that the results are thoroughly evaluated and validated across diverse LLM architectures and experimental conditions. Overall, we can see, the product $\mathbf{A}_{s}\times\mathbf{W}_{s}$ approximates the identity matrix, supporting the theoretical findings outlined in Corollary 4.2.

### v3

@@ I.1 LLMs Mimic the Human World Model, Not the World Itself
Interestingly, our texts reflect our mental activities and emotions, providing a window into this compressed world model. A recent study reveals a striking disparity between the limited information throughput of human behavior (approximately $10$ bits/s) and the vast sensory input available (around $10^{9}$ bits/s) [80]. This suggests that the human world model is an efficient, compressed abstraction, enabling us to reason, predict, and make decisions effectively. Despite this compression, humans have flourished as the dominant species on Earth, demonstrating the power of such a streamlined model.
@@
I.2 LLMs Versus Pure Vision Models: A Fundamental Difference
@@
Appendix J More Results on Llama-3 and DeepSeek-R1
We conduct additional experiments on recent LLMs, including Llama-3 and DeepSeek-R1, to further evaluate our findings. The experimental setup strictly adheres to the settings described in Section G.2. This enables a comprehensive investigation of our findings, ensuring that the results are thoroughly evaluated and validated across diverse LLM architectures and experimental conditions. Overall, we can see, the product $\mathbf{A}_{s}\times\mathbf{W}_{s}$ approximates the identity matrix, supporting the theoretical findings outlined in Corollary 4.3.
@@
Appendix K Future Directions
Embedding Causal Reasoning in LLMs Through Linear Unmixing
Our linear identifiability result lays a foundation for uncovering latent causal relationships among concepts, especially when these variables exhibit causal dependencies within the proposed latent variable model. By showing that the representations in LLMs are linear mixtures of latent causal variables, our analysis shows that linear unmixing of these representations may allow for the identification of underlying latent causal structures. This approach not only opens up the possibility of understanding causal dynamics within LLMs but also suggests that causal reasoning, particularly in latent spaces, could be achievable through the exploration of these linear unmixing techniques. We believe this work marks a pivotal step toward embedding robust causal reasoning capabilities into LLMs.